Write the slope-intercept form of the equation of the line through the given point and parallel to the given line. Point (15,1) Line 10x-2y=3
Answer by stanbon(75887)

Write the slope-intercept form of the equation of the line through the given point and parallel to the given line. Point (15,1) Line 10x-2y=3
---
Given line 10x-2y = 3
2y = 10x-3
y = 5x-3/2)
Slope = 5
------------
New equation with slope = 5 and point(15,1)
----
Form: y = mx + b
Solve for "b":
1 = 5*15+b
b = -74
----
Equation:
y = 5x-74
===============
cheers,
Stan H.
===============